Marcus Allen is evaluating a business opportunity to sell premium car wax at vintage car shows. The wax is sold in​64-ounce tubs. Marcus can buy the premium wax at a wholesale cost of $29 per tub. He plans to sell the premium wax for $84 per tub. He estimates fixed costs such as travel​ costs, booth rental​ cost, and lodging to be $880.

Requirement 1. Determine the number of tubs Marcus must sell per show to break even.

Begin by identifying the formula to compute the sales in units at various levels of operating income using the contribution margin approach.

(

+

) /

=

Sales in units

Marcus must sell

tubs per show to breakeven.

Requirement 2a. Determine the sales volume in units necessary to earn the desired profit assuming Marcus wants to earn a profit of 1,320 per show.

Marcus must sell

tubs per show to earn a profit of $1,320 per show.

b. Determine the sales volume in dollars necessary to earn the desired profit assuming Marcus wants to earn a profit of 1,320 per show. Begin by identifying the formula to compute the sales volume in dollars necessary to earn the desired profit.

x

=

Target sales in dollars

Marcus must achieve sales of $

c. Using the contribution margin​ format, prepare an income statement​ (condensed version) to confirm your answers to parts a and b.

Condensed Income Statement

Less:

Less:

Net income (loss)

Requirement 3. Determine the margin of safety between the sales volume at the breakeven point and the sales volume required to earn the desired profit. Determine the margin of safety in both sales​ dollars, units, and as a percentage.

Begin by identifying the formula to compute the margin of safety in dollars.

=

Margin of safety in dollars

The margin of safety in dollars is $

.

Now identify the formula to compute the margin of safety in units

=

Margin of safety in units

The margin of safety in units is

units.

Now identify the formula to compute the margin of safety as a percentage.

Margin of safety as a

/

=

percentage of target sales

The margin of safety as a percentage of target sales is

%.
Answer #1

(target profit+ fixed costs)/contribution Margin per unit = sales in units

= (0+880)/(84-29)

= 16

It must sell 16 tubes per show

2a required tubes = (1,320+880)/55

= 40 tubes

2b target sales = target sales in units*selling price per unit

= 40*84

= $3,360

Income statement

Sales 1,344 3,360
Less: variable cost 464 1,160
Contribution Margin 880 2,200
Less: fixed costs 880 880
Profit 0 1,320

Margin of safety in dollars = total sales - break even sales

Margin of safety = 3,360-1,344

= $2,016

% = Margin of safety sales/total sales

% = 2,016/3,360

= 60%